OpenStudy (anonymous):

Find the volume of this right circular cylinder. Dimensions are in centimeters. (Use 3.14 for π.) Radii=10 Diameter=20 Height=10

OpenStudy (paxpolaris):

\[\Large V=\pi \cdot r^2\cdot h\]